VEGAS Team Meeting

Attendees: Richard, Joe, Ray, Mark W., Dan, Glenn, Marty

Agenda:

General update

Meeting notes:

  1. The VEGAS hardware has been removed from the control room and located in the anechoic chamber. Tests are expected to begin Tuesday and will run between two and four days depending on the amount of mitigation that is required. When the system is returned to the control room, it will be fully configured with all connections in place for multi-bank testing.
  2. The schedule status circulated by Richard was agreed upon but there will be some re-ordering to accommodate sequencing discussions with the Ray & Joe. A revised version will be circulated later today (Monday).
  3. An email with the updated schedule will be sent to Lory and Rob along with a request to schedule the next PMO VEGAS meeting for the week of Jan 13.
  4. The participants agreed that there would be no further work on the ripples seen in Amanda's on-sky tests until we have some observational experience with VEGAS that would demonstrate a need for the issue to be addressed. If the filters have to be re-worked, it would be a major undertaking.

-- MartyBloss - 2014-01-06
Topic revision: r1 - 2014-01-06, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback